# Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication

### v2

@@ Abstract
By analyzing the spin-group symmetry of magnons, we establish two generalized Onsager’s relations in the magnon Hall effect, which reveals the rich and complicated structures of the magnon Berry curvature in the parameter space of different types of the exchange coupling. As an important consequence, we find that the isotropic exchange coupling is equally important as the Dzyaloshinskii-Moriya (DM) interaction in determining the magnon Hall effect, and it can transform to the latter under proper gauge transformation. Moreover, the diagonal and off-diagonal part of the symmetric anisotropic exchange coupling together can support the magnon Hall effect without the DM interaction. Both features are exemplified in ferromagnetic honeycomb lattices. Our work lays the ground for decoding the coupling between magnon transport and different types of exchange interactions.
@@
Introduction.—As potential carriers in energy and spin transport, magnons, elementary excitations of spin orders, have drawn much attention in recent years. One outstanding example is the magnon Hall effect [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17], which generates a thermal current perpendicular to the applied temperature gradient. The corresponding thermal conductivity can be related to the Berry curvature of magnon in the momentum space [5, 2, 1]. As a fundamental rule, the Onsager’s reciprocal relation[18, 19] restricts the structure of the transport coefficient. In the electronic counterpart, i.e., the anomalous Hall effect[20], it enforces the corresponding conductivity to be an odd function of the spin orientation, reveals the intimate relation between the band geometry and the order parameter, and hence serves as a guiding principle for experimental detection. In the magnonic case, however, besides separate case studies, a comprehensive understanding of the Onsager’s relation is still missing.
@@
The difficulty lies in the identification of the critical parameter in the magnon transport. It is widely believed that the Dzyaloshinskii-Moriya (DM) interaction[21, 22] serves as an order parameter, associated with the breaking of the effective time-reversal symmetry consisting of a spin rotation and real time-reversal operation [4, 23, 17], so that the thermal Hall conductivity is an odd function of the DM interaction. However, it has been long overlooked that due to the spin-group symmetry, the choice of the spin rotation axis contains large gauge freedom, which essentially mixes the isotropic, anisotropic and antisymmetric part of the exchanging coupling. This complexity has been implicitly envisioned in recent studies which show that the anisotropic exchange coupling [24, 25] may also induce a nontrivial magnon Hall effect [7]. To accurately identify the critical parameter in the magnon Hall effect, a detailed analysis of the effective time reversal symmetry from the spin-group perspective is required.
@@
In this work, we provide such a theory and derive two generalized Onsager’s relations. For any periodic magnetic textures, we use a local formulation of the magnonic Hamiltonian. We then identify the group $\tilde{G}$ of full effective time reversal operations (see Eq. (8)), inherited from the spin-only group[26, 27, 28, 29, 30, 31, 32, 33] in the original magnetic lattice. We find two types of effective time reversal operations, i.e., the uniform one and non-uniform one, from which two Onsager’s relations are established, revealing the rich and complicated dependence of the magnon Berry curvature on various exchange couplings. As important consequences, we find that the isotropic exchange coupling is as important as the DM interaction in determining the magnon Hall effect, and that the off-diagonal and diagonal part of the anisotropic exchange coupling can replace the role of the isotropic exchange coupling and DM interaction, respectively. Such findings are exemplified in magnons on ferromagnetic honeycomb lattices.
@@
Local formulation of magnons.—We first set up the general framework for discussing the effective time reversal operation and its implication. Without loss of generality, we assume that the spin texture still possesses the lattice translation symmetry and that there can be arbitrary number of spins in a unit cell, each of which can point along different directions in equilibrium. An example of such system is shown in Fig.1. Instead of choosing a global coordinate system for the spin texture, we adopt the following rule: we define local coordinate systems on different lattice sites such that the $z$-axis of the local frame always points along the equilibrium direction of the local spin order, as shown in Fig.1. Due to the periodicity of the spin texture, the local coordinate frame defined in this way is also periodic.
@@
Figure 1: Configuration of a general magnetic system. Different colors of spins represent different sublattices and spins belong to the same sublattice have same orientation and local coordinate.
@@
Here $\psi_{\bm{k}}=(a_{1,\bm{k}},\cdots,a_{N,\bm{k}},a^{\dagger}_{1,-\bm{k}},\cdots,a^{\dagger}_{N,-\bm{k}})^{T}$ represent the Nambu basis. The block of Eq. (2) can be expressed by $J^{ab}_{ni,mj}$ [35, 33, 14]
@@
where $\bm{R}_{ni}$ is the position of the $i$-th spin in the $n$-th unit cell, $J^{+}_{ni,mj}=J_{ni,mj}^{xx}+J_{ni,mj}^{yy}$, $J^{-}_{ni,mj}=J_{ni,mj}^{xx}-J_{ni,mj}^{yy}$, $D_{ni,mj}=J_{ni,mj}^{xy}-J_{ni,mj}^{yx}$, and $\Gamma_{ni,mj}=J_{ni,mj}^{xy}+J_{ni,mj}^{yx}$. By summing over the unit cell index, we can further define the following symbols according to Eq. (Generalized Onsager’s Relation in Magnon Hall Effect and Its Implication): $A_{ij}=J_{ij}^{+}-iD_{ij}-\delta_{ij}\sum_{mn\ell}J_{ni,m\ell}^{zz}$ and $B_{ij}=J_{ij}^{-}+i\Gamma_{ij}$. One can further check that $A(\bm{k})=A^{\dagger}(\bm{k})$ and $B(\bm{k})=B^{T}(-\bm{k})$. As a result, $\mathcal{H}$ is Hermitian.
Interestingly, the definition of the intermediate quantities $J^{\pm}$, $D$ and $\Gamma$ agrees with the decomposition of the exchange coefficient. With fixed lower indices, $J^{ab}_{ni,mj}$ is a rank-2 tensor, which can then be decomposed into three parts in the $xy$ subspace: the scalar part, also referred to as the isotropic exchange interaction, is represented by $J^{+}$, the antisymmetric part, also referred to as the DM interaction, is represented by $D$, the traceless symmetric part is represented by $J^{-}$ and $\Gamma$. We shall emphasize the $\pi/2$ phase difference between $J^{+}$ and $D$, and $J^{-}$ and $\Gamma$, and the fact that $J^{+}$ and $D$ work in the particle or hole sector but $J^{-}$ and $\Gamma$ mix particles with holes in the magnon Hamiltonian.
The Hamiltonian in Eq. (2) can be diagonalized by using para-unitary transformation[36, 37, 38, 39]
$T_{\bm{k}}^{\dagger}\mathcal{H}_{\bm{k}}T_{\bm{k}}=\begin{bmatrix}E_{\bm{k}}&\\ &E_{\bm{-k}}\end{bmatrix},$ (5)
where $E_{\bm{k}}=\mathrm{diag}(\varepsilon_{1,\bm{k}},\varepsilon_{2,\bm{k}},\cdots,\varepsilon_{N,\bm{k}})$. The structure of the eigen energy is due to particle-hole symmetry. In order to maintain bosonic commutation rules, the transformation matrix $T_{\bm{k}}$ has to satisfy a special orthonormal condition $T_{\bm{k}}^{\dagger}\Sigma_{z}T_{\bm{k}}=\Sigma_{z}$, where $\Sigma_{z}=\sigma_{z}\otimes\mathbb{I}_{N}$ with $\sigma_{z}$ being the Pauli matrix and $\mathbb{I}_{N}$ being the $N\times N$ unit matrix.
Within this framework, the thermal Hall conductivity of non-interacting magnons is given by [2, 5]
@@
where $c_{2}(\rho)=(1+\rho)(\mathrm{log}\frac{1+\rho}{\rho})^{2}-(\mathrm{log}\rho)^{2}-2\mathrm{Li}_{2}(-\rho)$ with $\mathrm{Li}_{2}(z)$ being the polylogarithm function, $\rho_{n}$ is Bose distribution function of the $n$-th band, and $\Omega_{n,\bm{k}}$ is Berry curvature [38]:
$(\Omega_{\mu\nu})_{n,\bm{k}}=i\left[\Sigma_{z}\frac{\partial T_{\bm{k}}^{\dagger}}{\partial k_{\mu}}\Sigma_{z}\frac{\partial T_{\bm{k}}}{\partial k_{\nu}}\right]_{nn}.$ (7)
Generalized Onsager’s Relation.—At the heart of the Onsager’s relation is the time reversal operation $T$. In the electronic Hall-type conductivity, on one hand, $T$ connects systems with opposite magnetic order; on the other hand, $T$ flips the sign of the Hall conductivity. The Onsager’s relation then states that the Hall conductivity is an odd function of the magnetic order[20, 40].
In the magnonic degree of freedom, however, the situation is quite different. Neither the Hamiltonian of the fluctuation in Eq. (1) nor the Berry curvature in Eq. (7) can be flipped by the time reversal operation $T$. Therefore, using $T$ alone cannot offer any useful information about the structure of the magnon Hall conductivity.
A practical generalization is the effective time reversal operation $\mathcal{T}=TC_{2x}^{s}$ [41, 42] where the rotation only acts on spin and flips the spin order. Since $\mathcal{T}$ can flip both the magnon Berry curvature and the DM interaction, it is usually assumed that the DM interaction plays the role of ‘magnetic order’ and is indispensable in the magnon Hall effect[3, 12, 4, 9, 10, 11, 13, 14, 16, 8, 15]. We shall emphasize that in previous study[23, 4, 17], this analysis is applied in ferromagnets and coplanar antiferromagnets where a global $C_{2}$ spin rotation axis exists.
In the general setting, a significant issue arises: what is the precise meaning of the $C_{2}$ spin rotation axis? For non-coplanar spin textures, there is no global spin rotation axis that can simultaneously flip all the spin orders. The best one can define is the local spin rotation axis that differ site by site, as employed in our framework and illustrated in Fig. 1.
The locally defined spin rotation axis makes prominent an elemental issue, i.e., the gauge freedom in choosing the local spin-x direction. At each site in a unit cell, only the spin-z direction is well defined, and any direction perpendicular to the $z$-th direction can work equally well as the $x$ direction. The spin-x axis on different sites can have no connection. This gauge freedom should also emerge in ferromagnetic and coplanar antiferromagnetic crystals.
@@
This group structure greatly restricts the magnon Hall effect. To see this, we make a gauge transformation that rotates the spin axis on the i-th site about the local $z$-axis by an angle $\phi_{i}$ with i going through all the local spin orders in a unit cell. As a result, we find that $A_{ij}\rightarrow A_{ij}e^{i(\phi_{i}-\phi_{j})}$ and $B_{ij}\rightarrow B_{ij}e^{i(\phi_{i}+\phi_{j})}$. Since the magnon Hall conductivity should not depend on any specific choice of $\phi_{i}$, the model parameters $J_{ni,mj}^{\pm}$, $D_{ni,mj}$ and $\Gamma_{ni,mj}$ should only appear in the form of gauge-independent clusters, the number of which is at least $N$-less than the total number of parameters in Eq. (2).
@@
We start with choosing $\mathcal{T}_{x}$ identically on each spin in a unit cell, which is the uniform effective time reversal operation. We emphasize that although the mathematical symbol is the same, $\mathcal{T}_{x}$ on different spins can involve $C_{2x}$ rotations along different axis in the lattice frame. In the magnon case, the effective time reversal operation $\mathcal{T}_{x}$ is the spinless-version of the time reversal operation, and can be identified as $\mathcal{T}_{x}=-K$ where $K$ stands for complex conjugation. We then find that[35] $\mathcal{T}_{x}\mathcal{H}_{\bm{k}}(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})\mathcal{T}_{x}^{-1}=\mathcal{H}_{-\bm{k}}^{*}(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})=\mathcal{H}_{\bm{k}}(J_{ij}^{+},J_{ij}^{-},-D_{ij},-\Gamma_{ij})$. The last equality holds by directly evaluating the Hamiltonian at $-\bm{k}$. By calculating the Berry curvature in the original parameter sets $(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})$ but with the time reversal operator and in the new parameter sets $(J_{ij}^{+},J_{ij}^{-},-D_{ij},-\Gamma_{ij})$, respectively, we find that
$(\Omega_{\mu\nu})_{n,\bm{k}}(J_{ij}^{\pm},D_{ij},\Gamma_{ij})=-(\Omega_{\mu\nu})_{n,-\bm{k}}(J_{ij}^{\pm},-D_{ij},-\Gamma_{ij}),$ (9)
We then obtain the first Onsager’s relation[35]:
@@
The effective time reversal symmetry can also be chosen in a non-uniform manner. We classify the spins in a unit cell into two groups, such that the time reversal operation on spins in the first group is $\mathcal{T}_{x}$ while that on spins in the remaining group is $\mathcal{T}_{y}$. There are consequently two types of parameters $J_{ij}^{\pm}$, $D_{ij}$, and $\Gamma_{ij}$, i.e., those with $i$ and $j$ within the same group, labeled by $J_{ij}^{\pm,intra}$, $D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}$ and those with $i$ and $j$ belonging to different groups, labeled by $J_{ij}^{\pm,inter}$, $D_{ij}^{inter}$, and $\Gamma_{ij}^{inter}$. Since $\mathcal{T}_{y}=K$, we then find that the total effective time reversal symmetry reads as $\mathcal{T}=(-KI_{n_{1}})\oplus(KI_{n_{2}})$ with $n_{1}$ and $n_{2}$ being the number of spins in the first and second group, respectively. The effect of $\mathcal{T}$ on the intra-group parameters is the same as in the uniform case, i.e., $J_{ij}^{\pm,intra}\rightarrow J_{ij}^{\pm,intra}$, $D_{ij}^{intra}\rightarrow-D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}\rightarrow-\Gamma_{ij}^{intra}$. Due to the relative minus sign between $\mathcal{T}_{x}$ and $\mathcal{T}_{y}$ in $\mathcal{T}$, the effect on the inter-group parameters is opposite: $J_{ij}^{\pm,intra}\rightarrow-J_{ij}^{\pm,intra}$, $D_{ij}^{intra}\rightarrow D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}\rightarrow\Gamma_{ij}^{intra}$. We can then obtain the second Onsager’s relation[35]:

### v3

# Magnon Thermal Hall Effect Induced By Symmetric Exchange Interaction

@@ Abstract
By analyzing the spin-group symmetry of magnons, we establish two generalized Onsager’s relations in the magnon thermal Hall effect, which reveals the rich and complicated structures of the magnon Berry curvature in the parameter space of different types of the exchange coupling. As an important consequence, we find that the diagonal and off-diagonal part of the symmetric anisotropic exchange coupling together can support the planar magnon thermal Hall effect without the DM interaction. This removes the local intersion-symmetry-breaking condition for generating the magnon thermal Hall effect. Moreover, we predict an exotic phenomenon that the thermal conductivity exhibits angular dependence with respect to the in-plane magnetization. Our work lays the ground for decoding the coupling between magnon transport and different types of exchange interactions.
@@
Introduction.—As potential carriers in energy and spin transport, magnons, elementary excitations of spin orders, have drawn much attention in recent years. One outstanding example is the magnon thermal Hall effect[1], which generates a thermal current perpendicular to the applied temperature gradient. The corresponding thermal conductivity can be related to the Berry curvature of magnon in the momentum space [2, 3, 4, 5]. As a fundamental rule, the Onsager’s reciprocal relation[6, 7] restricts the structure of the transport coefficient. In the electronic counterpart, i.e., the anomalous Hall effect[8], it enforces the corresponding conductivity to be an odd function of the spin orientation, reveals the intimate relation between the band geometry and the order parameter, and hence serves as a guiding principle for experimental detection. In the magnonic case, however, besides separate case studies, a comprehensive understanding of the Onsager’s relation is still missing.
@@
The difficulty lies in the identification of the critical parameter in the magnon transport. It is widely believed that the Dzyaloshinskii-Moriya (DM) interaction[9, 10] is indispensable in the magnon thermal Hall effect, associated with the breaking of the effective time-reversal symmetry consisting of a spin rotation and real time-reversal operation [11, 12, 13, 14, 15], so that the thermal Hall conductivity is an odd function of the DM interaction. Magnon thermal Hall effect induced by DM interaction has been studied in various systems theoretically and experimentally, such as honeycomb lattice[16, 17, 18], kagome lattice[19, 20, 13, 21, 22, 23, 24], square lattice[25, 26] and Lieb lattice[27]. However, it has been long overlooked that due to the spin-group symmetry, the choice of the spin rotation axis contains large gauge freedom, which essentially mixes the isotropic, anisotropic and antisymmetric part of the exchanging coupling. This complexity has been implicitly envisioned in recent studies [28] which show that the anisotropic exchange coupling [29, 30] which comes from second order perturbation of spin-orbit coupling may also induce a nontrivial magnon thermal Hall effect. To accurately identify the critical parameter in the magnon thermal Hall effect, a detailed analysis of the effective time reversal symmetry from the spin-group perspective is required.
@@
In this work, we provide such a theory and derive two generalized Onsager’s relations, by identifying the group $\tilde{G}$ of full effective time reversal operations (see Eq. (6)), inherited from the spin-only group[31, 32, 33, 34, 35, 36, 37, 38] in the original magnetic lattice, and recognizing two types of effective time reversal operations. They reveal the rich and complicated dependence of the magnon Berry curvature on various exchange couplings. As an important consequence, we establish the necessary condition when the anisotropic interaction alone can induce magnon thermal Hall without the DM interaction. Since the anisotropic exchange coupling does not need to break the local inversion symmetry, our theory then expands potential systems with the magnon thermal Hall effect. Examples are $\rm VAu_{4}$} and monolayer $\rm CrCl_{3}$}. Moreover, we show that the anisotropic exchange coupling can lead to an exotic in-plane magnon thermal Hall effect where the thermal Hall conductivity depends on in-plane orientation of magnetization.
@@
Local formulation of magnons.—We first set up the general framework for discussing the effective time reversal operation and its implication. We consider periodic spin textures and assume that there can be arbitrary number of spins in a unit cell, each of which can point along different directions. Instead of choosing a global coordinate system for the spin texture, we adopt the following rule: we define local coordinate systems on different lattice sites such that the $z$-axis of the local frame always points along the equilibrium direction of the local spin order. Such local coordinate frame defined in this way is periodic.
@@
Here $\psi_{\bm{k}}=(a_{1,\bm{k}},\cdots,a_{N,\bm{k}},a^{\dagger}_{1,-\bm{k}},\cdots,a^{\dagger}_{N,-\bm{k}})^{T}$ represent the Nambu basis. The block of Eq. (2) can be expressed by $J^{ab}_{ni,mj}$ [40, 38]
@@
where $\bm{R}_{ni}$ is the position of the $i$-th spin in the $n$-th unit cell, $J^{+}_{ni,mj}=J_{ni,mj}^{xx}+J_{ni,mj}^{yy}$, $J^{-}_{ni,mj}=J_{ni,mj}^{xx}-J_{ni,mj}^{yy}$, $D_{ni,mj}=J_{ni,mj}^{xy}-J_{ni,mj}^{yx}$, and $\Gamma_{ni,mj}=J_{ni,mj}^{xy}+J_{ni,mj}^{yx}$. By summing over the unit cell index, we can further define the following symbols according to Eq. (Magnon Thermal Hall Effect Induced By Symmetric Exchange Interaction): $A_{ij}=J_{ij}^{+}-iD_{ij}-\delta_{ij}\sum_{mn\ell}J_{ni,m\ell}^{zz}$ and $B_{ij}=J_{ij}^{-}+i\Gamma_{ij}$. One can further check that $A(\bm{k})=A^{\dagger}(\bm{k})$ and $B(\bm{k})=B^{T}(-\bm{k})$. As a result, $\mathcal{H}$ is Hermitian.
Interestingly, the definition of the intermediate quantities $J^{\pm}$, $D$ and $\Gamma$ agrees with the decomposition of the exchange coefficient. With fixed lower indices, $J^{ab}_{ni,mj}$ is a rank-2 tensor, which can then be decomposed into three parts in the $xy$ subspace: the scalar part, also referred to as the isotropic exchange interaction, is represented by $J^{+}$, the antisymmetric part, also referred to as the DM interaction, is represented by $D$, the traceless symmetric part is represented by $J^{-}$ and $\Gamma$, which originates from the spin-orbit coupling just as the DM interaction. We shall emphasize there is $\pi/2$ phase difference between $J^{+}$ and $D$, as well as $J^{-}$ and $\Gamma$.
Finally, the thermal Hall conductivity of non-interacting magnons is given by [3, 2]
@@
where $c_{2}(\rho)=(1+\rho)(\mathrm{log}\frac{1+\rho}{\rho})^{2}-(\mathrm{log}\rho)^{2}-2\mathrm{Li}_{2}(-\rho)$ with $\mathrm{Li}_{2}(z)$ being the polylogarithm function, $\rho_{n}$ is Bose distribution function of the $n$-th band, and $\Omega_{n,\bm{k}}$ is the magnon Berry curvature [41]. We ignore the magnon-magnon interaction in the magnetic Hamiltonian which only causes extrinsic contribution to the magnon thermal Hall effect and can be disentangled from the intrinsic one by studying temperature profile.
Before closing this section, we want to comment on the role of magnetic field. The equilibrium magnetic state often depends on external magnetic field. As for its fluctuation, the effect of magnetic field is more involving. In collinear ferromagnets, the magnetic field does not explicitly enter the magnon Hamiltonian in Eq. (1). Yet, it can still affects the magnon thermal Hall effect by re-orienting the equilibrium direction of the spin order. For nonuniform spin orders such as noncollinear antiferromagnetism, the magnetic field can enter Eq. (1) through the Zeeman term: $B\sum_{ni}\cos(\theta_{i})S_{ni}^{z}$, where $\theta_{i}$ is the angle between the magnetic field and the local spin order. It can be put in a bilinear form of $S_{ni}^{x}$ and $S_{ni}^{y}$. In this way it can affect the magnon thermal Hall effect but it does not change the role of the symmetric exchange interaction.
Generalized Onsager’s Relation.—At the heart of the Onsager’s relation is the time reversal operation $T$. In the electronic anomalous Hall effect, $T$ can flip both the spin order and the anomalous Hall conductivity. The Onsager’s relation then states that the anomalous Hall conductivity is an odd function of the magnetic order [8, 42]. In the magnonic degree of freedom, however, neither the magnon Hamiltonian in Eq. (1) nor the Berry curvature can be flipped by $T$. Therefore, using $T$ alone cannot offer any useful information about the structure of the magnon thermal Hall conductivity.
A practical generalization is the effective time reversal operation $\mathcal{T}=TC_{2x}^{s}$ [11, 12, 13, 15, 14] where the rotation only acts on spin. Since $\mathcal{T}$ can flip both the magnon Berry curvature and the DM interaction, it is usually assumed that the DM interaction parameter plays the role of ‘magnetic order’ and is hence indispensable in the magnon thermal Hall effect[1, 20, 13, 16, 17, 27, 21, 26, 19, 18]. This analysis has been applied in ferromagnets and coplanar antiferromagnets where a global $C_{2}$ spin rotation axis exists[14, 13, 15].
However, for general periodic spin orders, there is a gauge freedom in choosing the local spin-x direction[40]. At each site in a unit cell, only the spin-z direction is well defined, and any direction perpendicular to the $z$-th direction can work equally well as the $x$ direction. The spin-x axis on different sites can have no connection. This gauge freedom also exists in ferromagnetic and coplanar antiferromagnetic crystals.
@@
We start with choosing $\mathcal{T}_{x}$ identically on each spin in a unit cell, which is the uniform effective time reversal operation. We emphasize that although the mathematical symbol is the same, $\mathcal{T}_{x}$ on different spins can involve $C_{2x}$ rotations along different axis in the lattice frame. In the magnon case, the effective time reversal operation $\mathcal{T}_{x}$ is the spinless-version of the time reversal operation, and can be identified as $\mathcal{T}_{x}=-K$ where $K$ stands for complex conjugation. We then find that[40] $\mathcal{T}_{x}\mathcal{H}_{\bm{k}}(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})\mathcal{T}_{x}^{-1}=\mathcal{H}_{-\bm{k}}^{*}(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})=\mathcal{H}_{\bm{k}}(J_{ij}^{+},J_{ij}^{-},-D_{ij},-\Gamma_{ij})$. The last equality holds by directly evaluating the Hamiltonian at $\bm{k}$. By calculating the Berry curvature in the original parameter sets $(J_{ij}^{+},J_{ij}^{-},D_{ij},\Gamma_{ij})$ but with the time reversal operator and in the new parameter sets $(J_{ij}^{+},J_{ij}^{-},-D_{ij},-\Gamma_{ij})$, respectively. We then obtain the first Onsager’s relation[40]:
@@
The effective time reversal symmetry can also be chosen in a non-uniform manner. We classify the spins in a unit cell into two groups, such that the time reversal operation on spins in the first group is $\mathcal{T}_{x}$ while that on spins in the remaining group is $\mathcal{T}_{y}$. There are consequently two types of parameters, i.e., those with $i$ and $j$ within the same group, labeled by $J_{ij}^{\pm,intra}$, $D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}$ and those with $i$ and $j$ belonging to different groups, labeled by $J_{ij}^{\pm,inter}$, $D_{ij}^{inter}$, and $\Gamma_{ij}^{inter}$. The effect of $\mathcal{T}$ on the intra-group parameters is the same as in the uniform case, i.e., $J_{ij}^{\pm,intra}\rightarrow J_{ij}^{\pm,intra}$, $D_{ij}^{intra}\rightarrow-D_{ij}^{intra}$, and $\Gamma_{ij}^{intra}\rightarrow-\Gamma_{ij}^{intra}$. In contrast, its effect on the inter-group parameters is opposite: $J_{ij}^{\pm,inter}\rightarrow-J_{ij}^{\pm,inter}$, $D_{ij}^{inter}\rightarrow D_{ij}^{inter}$, and $\Gamma_{ij}^{inter}\rightarrow\Gamma_{ij}^{inter}$. We can then obtain the second Onsager’s relation[40]: